A motorboat at full throttle takes two hours
longer to travel 75 miles against the current than it takes to
travel the same distance with the current. If the rate of the
current is 5 mph, then what is the speed of the boat at full
throttle in still water?
I have been working on this problem and I can't figure it out. I have to provide detailed soultions step by step

Answer by scott8148(6628)

time = distance / rate ___ there is a 2 hr difference in the times

75 / (s - 5) = [75 / (s + 5)] + 2

multiplying by (s-5)(s+5) ___ 75(s+5) = 75(s-5) + 2(s-5)(s+5) ___ 75s + 375 = 75s - 375 + 2s^2 - 50

800 = 2s^2